OpenStudy (anonymous):

what happens if you get something like this? 15=22?

OpenStudy (anonymous):

5x+15-3x=22+22

OpenStudy (♪chibiterasu):

5x + 15 - 3x = 22 + 22 Is this the original equation given?

OpenStudy (♪chibiterasu):

We are trying to find x, we need to isolate x.

OpenStudy (♪chibiterasu):

What steps did you take to solve this?

OpenStudy (christos):

5x-3x = ?